OpenStudy (anonymous):

Alicia can row 7 miles downstream in the same time it takes her to row 4 miles upstream. She rows downstream 5 miles/hour faster than she rows upstream. Find Alicia’s rowing rate each way.

OpenStudy (sriram):

let x be the rate downstream & y be the speed added by river x+y= 7/t x-5-y=4/t as both 't' r equal

OpenStudy (sriram):

stuck here!!

OpenStudy (sriram):

7/x+y = 4 / x-5-y cross multiply 7x-45-7y=4x+4y 3x-11y=45 x>5 and x-y-5>0 x>y+5 i guess we may get more than 1 ans. for this

OpenStudy (a_clan):

let speed of boat =Vb speed of river =Vr Given, Vr+Vb = 7/t (1) ----speed of rowing downstream(i.e. speed of boat alongwith river) Vr-Vb = 4/t (2)----speed of rowing upstream(i.e. speed of boat against the flow of river) Also given, (Vr+Vb) - (Vr-Vb) = 5 mph 2Vb = 5 Vb = 5/2 mph Putting back in eqn (1) and (2), we get Vr + 5/2 = 7/t and Vr - 5/2 = 4/t On, solving these two equations simultaneously, we get t= 3/5 hr So, rowing downstream rate = 7/t = 35/3 mph rowing upstream rate = 4/t = 20/3 mph